OpenStudy (anonymous):

A study of five hundred adults found that the number of hours they spend on social networking sites each week is normally distributed with a mean of 14 hours. The population standard deviation is 3 hours. What is the margin of error for a 95% confidence interval? 0.134 0.220 0.263 0.313

ganeshie8 (ganeshie8):

Margin of error for 95% CI = \(\large 1.96 \times \frac{\sigma}{\sqrt{n}}\)

ganeshie8 (ganeshie8):

plugin given values and simplifyy